OpenStudy (anonymous):

A quadratic equation is shown below: x2 + 18x + 76 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)2 = q, where p and q are integers? A. Add 9 to both sides of the equation B Add 5 to both sides of the equation C. Subtract 5 from both sides of the equation <---- My answer D. Subtract 9 from both sides of the equation

Directrix (directrix):

I think this question is about writing the left side of the quadratic as a perfect square trinomial.

Directrix (directrix):

x2 + 18x + 76 = 0 Take half the coefficient of the x term, 18, and then square it. 1/2*18 = 9. 9^2 = 81

Directrix (directrix):

x2 + 18x + 18 = -76 + 18

Directrix (directrix):

Why did you subtract 5 - just wondering.

OpenStudy (anonymous):

oh ok I see what you mean (:

Directrix (directrix):

My technique is not working. Let's test the options.

OpenStudy (anonymous):

ok

Directrix (directrix):

form (x - p)2 = q, where p and q are integers I will try adding 5 to both sides x2 + 18x + 76 = 0 x2 + 18x + 76 + 5 = 0 + 5 x^2 + 18x + 81 = 5 (x + 9)^2 = 5

OpenStudy (anonymous):

oh ok

Directrix (directrix):

Yes, but there is a plus where there should be a minus. p and q are integers so the expression could be written as: (x + (-9))^2 = 5

Directrix (directrix):

I don't know about that. Let's try what you think is correct.

OpenStudy (anonymous):

ok (:

Directrix (directrix):

x2 + 18x + 76 = 0 x2 + 18x + 76 -5 = 0 -5 x^2 + 18x + 71 = -5 This --> x^2 + 18x + 71 is not a perfect square trinomial so that is not the answer.

OpenStudy (anonymous):

oh ok I had a feeling at first that it wasn't

Directrix (directrix):

Whatever number is added or subtracted has to turn 76 into a a perfect square.

Directrix (directrix):

The only option that does that is this: B Add 5 to both sides of the equation provided that (x + [-9] )^2 = 5 is an acceptable form for the answer. I don't see why not because of this: form (x - p)2 = q, where p and q are integers

OpenStudy (anonymous):

oh ok yea I agree with completely (;

Directrix (directrix):

I was just looking at some similar problems and think that the B Add 5 is correct.

OpenStudy (anonymous):

oh ok thank you so much again (:, is it ok if u could help with a few more question if that's ok ?

Directrix (directrix):

Okay, post in a new thread and close this one, okay?

OpenStudy (anonymous):

ok (: